The consumption function is given by the equation: C = 400 + 0.8(Yd) where C =...

  1. The consumption function is given by the equation: C = 400 + 0.8(Yd) where C = consumption and Yd = disposable income. Currently, the level of disposable income is $5000.
    1. What is the current level of consumption?
    2. What is the current level of saving?
    3. What is the MPC? What is the MPS?
    4. Calculate the APC and the APS.

6. We are given the following information about the levels of disposable income and savings:

Disposable Income

Consumption

Savings

APC

APS

MPC

MPS

$2400

$-80

---

---

2600

-40

2800

0

3000

40

3200

80

3400

120

3600

160

3800

200

4000

240

  1. Complete the table.
  2. What is the break-even level of income?
Answer #1

Q) C = 400 + 0.8(Yd)

Yd= $5000

a) C = 400 + 0.8(Yd)

Putting the value of disposable income in the above equation,

C = 400 + 0.8(5000)

= 400 + 4000

= $4400 is the answer.

So, the current level of consumption is $4400.

b) Savings = Disposable income - Consumption

= 5000 - 4400

= $600 is the answer.

So, the current level of savings is $600.

c) C = 400 + 0.8(Yd)

MPC is calculated by differentiating the consumption function with respect to disposable income as it measures the change in consumption due to a given change in income.

So, differentiating the consumption function with respect to disposable income,

MPC = dC / dYd = C / Yd = 0.8 is the answer.

MPC + MPS = 1

0.8 + MPS = 1

MPS = 1 - 0.8 = 0.2 is the answer.

d) APC = consumption / disposable income

= 4400 / 5000

= 0.88 is the answer.

APC + APS = 1

0.88 + APS = 1

APS = 1 - 0.88 = 0.12 is the answer.